### v1

# AI-Generated Images Introduce Invisible Relevance Bias to Text-Image Retrieval

Abstract.
With the advancement of generation models, AI-generated content (AIGC) is becoming more realistic and less discernible, flooding the Internet. A recent study suggests that this phenomenon has elevated the issue of source bias in text retrieval for web searches. Specifically, neural retrieval models tend to rank generated texts higher than human-written texts.
In this paper, we extend the study of this bias to cross-modal retrieval. Firstly, we successfully construct a suitable benchmark to explore the existence of the bias. In this benchmark, AI-generated images possess visual semantics sufficiently similar to real images. Subsequent extensive experiments on this benchmark reveal that AI-generated images introduce an invisible relevance bias to text-image retrieval models. Specifically, our experiments show that text-image retrieval models tend to rank the AI-generated images higher than the real images, even though the AI-generated images do not exhibit more visually relevant features to the query than real images. This invisible relevance bias is prevalent across retrieval models with varying training data and architectures, encompassing models trained from scratch to those pre-trained on extensive image-text pairs, and spanning dual-encoder retrieval models to fusion-encoder retrieval models. Furthermore, our subsequent exploration reveals that the inclusion of AI-generated images in the training data of the retrieval models exacerbates the invisible relevance bias. The above phenomenon triggers a vicious cycle in which AI-generated images have a higher chance of being obtained from massive amounts of data, which makes them more likely to be mixed into the training of retrieval models and such training makes the invisible relevance bias become more and more serious.
To elucidate the potential causes of invisible relevance and address the aforementioned issues, we introduce an effective training method aimed at alleviating the invisible relevance bias. Subsequently, we apply our proposed debiasing method to retroactively identify the causes of invisible relevance, revealing that the AI-generated images induce the image encoder to embed additional information into their representation. This information exhibits a certain consistency across generated images with different semantics and can make the retriever estimate a higher relevance score. We conduct experiments to substantiate this assertion.
1. Introduction
With the advancement of generation models, AI-generated content (AIGC) has become increasingly realistic and less distinguishable (Yang et al., 2023; Brown et al., 2020). The utilization of AI for content generation has significantly transformed the landscape of content creation. It has not only reduced the cost of content generation but has also enhanced efficiency, leading to a rapid influx of a substantial volume of AI-generated content onto the Internet (Ai et al., 2023; Dai et al., 2023).
Figure 1. Bias found in our paper. IR models tend to rank AI-generated images higher than real images even though they have very similar semantics. This bias increases the likelihood of the generated images being extensively exposed within vast datasets, which makes them more likely to be mixed into the training of AIGC and retrieval models, leading to more serious bias and forming a vicious cycle.
Information retrieval (IR) is an important way for people to obtain the target information from massive Internet information (Manning, 2009). However, the rapid proliferation of AI-generated content (AIGC) presents a significant new challenge to IR. Specifically, as the Internet becomes increasingly populated with AI-generated content, the question arises: How will these contents influence the ranking results of retrieval systems? In response to this challenge, (Dai et al., 2023) conducted a study on text modality and found that neural information retrieval models exhibit a preference for texts generated by Large Language Models (LLMs). They refer to this category of biases in neural retrieval models towards the LLM-generated text as the source bias. However, beyond textual information, the Internet is replete with a substantial number of images, serving as crucial sources for IR systems. Building on this, our paper extends the investigation of source bias in AI-generated content to text-image retrieval models. A pivotal question emerges: What impact will AI-generated images have on existing text-image retrieval models?
A crucial distinction from research on text modality lies in the methodology for constructing a benchmark to simulate retrieval scenarios comprising both real images and AI-generated images for IR models. A reasonable scenario for assessing the potential bias requires that the generated images and the real images have sufficiently similar visual semantics. This can avoid increasing or decreasing some semantic associations between the generated images and the query caused by the introduction or lack of some semantic information during the image generation. In other words, the IR model preferring (or rejecting) an AI-generated image that is more (or less) semantically similar to the query than the real image cannot conclusively prove the existence (or nonexistence) of the bias. Study on text modality prompts LLMs to rewrite each real text in the corpus, thereby constructing an LLM-generated text for each real text without introducing additional semantic information. However, this approach is unsuitable for constructing the data in image modality. It is because we find mainstream image generation models such as diffusion models (Rombach et al., 2022) are not good at following prompts to rewrite images while maintaining semantics. Besides, the rewriting paradigm diverges from the prevalent content generation scenario, where receiving a description as a prompt and generating the texts or images. To solve the above problem, we propose an image over-sampling and selection strategy based on the generation according to the merged caption. Experimental results and human evaluation show that our proposed method can successfully construct semantically similar AI-generated images for real images. The mixture of these images provides an effective simulation and validation benchmark for investigating text-image retrieval models within scenarios featuring a mix of AI-generated and real images. Our experiments on this benchmark reveal that text-image retrieval models tend to rank the AI-generated images higher than the real images, even though the AI-generated images do not exhibit more visually relevant features to the query than real images. We define this bias as invisible relevance bias introduced by AI-generated images.
Another important point that is not considered in previous work is that AI-generated content does not only have an impact on the inference stage of the retrieval models. Due to their large-scale distribution on the Internet, they are very likely to be mixed into the training data of retrieval models in the future. Therefore, we further mix the AI-generated images into the training of the retrieval model and explore the rules of its impact on the retrieval models by adjusting the mixing ratio. Our experimental results show that as the mixing ratio of AI-generated images increases, the invisible relevance bias becomes more serious, and the retrieval performance of on the generated images gradually becomes stronger, while the retrieval performance on the real images is gradually damaged. This phenomenon reveals that invisible relevance bias causes the generated images to have a higher chance of being obtained from massive data, which makes them more likely to be mixed into the training of AIGC and retrieval models (Shumailov et al., 2023), leading to more serious bias and forming a vicious cycle. (shown in Figure 1)
To mitigate the bias, we propose an effective training method for debiasing. Specifically, we introduce a contrastive loss between real and AI-generated images during training. This contrastive loss measures the invisible relevance bias introduced by the AI-generated images for retrieval models. Optimization of this loss can reduce the invisible relevance bias, thereby alleviating the preference for AI-generated images in text-image retrieval models. In addition, we also introduce the sampling probability to enable our debiasing method to dynamically adjust the preference to AI-generated images of text-image retrieval models.
Last but not least, we use our proposed debiasing method to reversely determine that the causes of invisible relevance is the generated images causing the image encoder to embed additional information into their representation. This information has a certain consistency in different generated images and can make the retriever estimate a higher relevance score. We also design experiments to support this.
The contributions of this paper can be concluded as:
(1) We reveal the potential impact of AI-generated images on the ranking results of text-image retrieval systems in the context of the rapid development of AIGC. To reasonably study the impact, we propose an effective method to construct a reasonable benchmark to simulate text-image retrieval in scenarios consisting of real and AI-generated images.
(2) Experimental results show that AI-generated images introduce the invisible relevance to text-image retrieval models, causing the bias that retrieval models prefer ranking AI-generated images higher even though the AI-generated images do not exhibit more visually relevant features to the query than real images. This bias widely exists in retrieval models with different training data and architectures, from models trained from scratch to models pre-trained on massive image-text pairs, from dual-encoder retrieval models to fusion-encoder retrieval models. We also reveal that the loop formed by training and inference causes the retrieval model to fall into a vicious cycle of increasingly serious bias.
(3) We propose an effective method to mitigate the bias by measuring and optimizing the invisible relevance bias introduced by AI-generated images in the training of retrieval models.
(4) We determine the causes of invisible relevance bias is that AI-generated images make the image encoder in the retriever embed additional information to their representations. This additional information is invisible but can amplify the relevant information to get a higher relevance score.
2. Benchmark Construction
This paper focuses on revealing the potential impact of AI-generated images on the ranking results of text-image retrieval systems. The prerequisite for all this is to construct a retrieval database consisting of real images and AI-generated images. This database should meet four critical requirements. (-$\mathcal{R}\mbox{-}1$) Firstly, from the perspective of fair assessment, a reasonable scenario for assessing the potential bias requires that the generated images and the real images must have sufficiently similar visual semantics. This can avoid increasing or decreasing some semantic associations between the generated images and the query caused by the introduction or lack of some semantic information during the image generation. That is, the IR model preferring (or rejecting) an AI-generated image that is more (or less) semantically similar to the query than the real image cannot prove the existence (or nonexistence) of the bias. (-$\mathcal{R}\mbox{-}2$) Secondly, retrieval performance on the database containing only generated images should not change significantly compared to retrieval performance on real images only. This can further ensure that no additional visual semantics relevant (or irrelevant) to the query are introduced during the image generation. (-$\mathcal{R}\mbox{-}3$) Thirdly, the image generation pipeline should be in line with the most common scenario of content generation that receives a description as prompt and generates the texts or images, so that the obtained bias assessment results can be more consistent with the actual scene. (-$\mathcal{R}\mbox{-}4$) Fourthly, the number of generated images and real images should be equal, so as to avoid additional bias caused by the difference in the number of images. Based on the above three requirements, we propose an image over-sampling strategy based on the generation via merged caption, which will be introduced in the following contents.
2.1. Image Generation
We select two widely used text-image retrieval datasets including Flickr30k (Plummer et al., 2015) and MSCOCO (Lin et al., 2014) as the basis for our benchmark construction. In these two datasets, each image is annotated with five captions that describe the content of the image. For each real image, we aim to generate a corresponding AI-generated image. We propose four image generation methods and use the above four requirements to select the optimal image generation method. Details will be introduced in the following.
2.1.1. Generation Based on Single Caption
In this method, we randomly select one caption from the five captions of each real image and use the selected caption as the prompt of the stable diffusion model [^1] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 to obtain the AI-generated image corresponding to the real image.
2.1.2. Generation Based on Merged Caption
Since a single caption is not guaranteed to contain the complete visual semantics of the image, we propose to merge five captions to obtain a relatively complete semantic description of the image. Specifically, for each real image ($I_{r}$), we use gpt-3.5-turbo to combine the five captions to get a new merged caption. And input the merged caption to a stable diffusion model $\mathbf{M}$ to generate the image ($I_{g}$), which can be described as:
[…]
$I^{g}$ $=\mathbf{M}(C_{m}).$
The intention of this design is that different captions of an image may be the descriptions from different perspectives of the image. Therefore, merging these captions can obtain an overall description of the image, thereby making stable diffusion generate images that have sufficiently similar visual semantics to the real image as much as possible. Besides, this also avoids the relevance bias in assessment caused by the generated image being overly inclined to one certain caption.
2.1.3. Generation Based on Image Rewriting
We not only input the merged caption $C_{m}$ but also input the real image $I_{r}$ as the initial image to condition the generation of the new image. Generation constrained by the input real image can output the AI-generated image that has higher similarity with the real image.
Table 1. Comparison between different generation methods under four requirements mentioned at the beginning of Section 2. Similarity is the cosine between embeddings of real and generated images encoded by OpenClip. Retrieval Performance is the difference in retrieval performance of BEIT-3 on the databases only containing generated images and only containing real images respectively.
Similarity (-$\mathcal{R}\mbox{-}1$) Retrieval Performance (-$\mathcal{R}\mbox{-}2$) Generation Pipeline (-$\mathcal{R}\mbox{-}3$) Ratio of Number (-$\mathcal{R}\mbox{-}4$)
Single Caption 0.5275 $\left|\triangle\rm{NDCG}@5\right|=10.8$ text to image ($\checkmark$) 1:1 ($\checkmark$)
[…]
2.1.4. Generation Based on Image Over Sampling and Selection
In order to further narrow the semantic similarity between the generated images and real images, we propose an image over-sampling strategy. Specifically, for a real image $I^{r}$, we use stable diffusion to perform multiple times generation with different random seeds and get $n$ generated images $\{{I^{g}_{1}},{I^{g}_{2}},...,{I^{g}_{n}}\}$. Then, we use the vision encoder $v(\cdot;\theta)$ of a powerful open source pre-trained vision-language model [^2] https://huggingface.co/laion/CLIP-ViT-H-14-laion2B-s32B-b79K to get the embedding $\bm{e_{r}}$ for $I^{r}$ and set of embeddings $\bm{E}=\{\bm{{e^{g}_{1}}},\bm{{e^{g}_{n}}},...,\bm{{e^{g}_{n}}}\}$ for $\{{I^{g}_{1}},{I^{g}_{2}},...,{I^{g}_{n}}\}$. We calculate the cosine similarity between each embedding in $\bm{E}$ and $\bm{e_{r}}$ to get the similarity between the generated images and the real image $I^{r}$. Finally, we select the generated image with the maximum similarity between $I^{r}$ as the final generated image ${I}^{g}$ for the real image:
(1) $I^{g}=\arg\max_{i}\{dot(\bm{e^{r}},\bm{e^{g}_{i}})|\bm{{e^{g}_{i}}}\in\bm{E}\}.$
2.1.5. Comparison between Different Generation Methods
The comparison between different generation methods under the above-mentioned four requirements is shown in Table 1. Taking these results into account, we choose generation based on merged caption and image over-sampling and selection as the image generation method (shown in Figure 2) to construct our benchmark. Using this method, we can generate the images for each real image in Flicker30k and MSCOCO. For the convenience of description, in the following contents, Flicker30k and MSCOCO indicate the datasets that only contain real images, Flicker30k (AIGC) and MSCOCO (AIGC) indicate the datasets that only contain AI-generated images, Flicker30k+AIGC and MSCOCO+AIGC indicate the datasets that consist of both real and AI-generated images, each real image has an AI-generated image.
Figure 2. Pipeline of our benchmark construction.
2.2. Assessing the Quality of Generated Images
Generated images and real images having sufficiently similar visual semantics so as to avoid introducing more semantics relevant to the query is a prerequisite for fairly assessing the bias caused by invisible relevance. In this section, we propose two methods to measure the semantic similarity. Firstly, we perform retrieval on two databases that only contain real images and only contain generated images respectively, and count the retrieval performance. The intention for this is that if the generated images have more semantics relevant to their corresponding queries than the real images, images in the database will become easier to distinguish, resulting in the retrieval performance on the database that only contains generated images significantly higher than the performance on the database that only contains real images. Secondly, we introduce human evaluation to further evaluate the semantic information from human vision. Specifically, we let humans determine whether the generated images have more visual semantics relevant to the queries compared to the real images.
2.2.1. Retrieval Performance
Table 2. Performance of the retrieval models on the databases that only contain real images and only include AI-generated images respectively. Performance on the AI-generated images is not significantly better than the performance on real images can demonstrate the success of our benchmark construction strategy.
Flicker30k or Flicker30k (AIGC) MSCOCO or MSCOCO (AIGC)
NDCG@1 NDCG@3 NDCG@5 R@1 R@3 R@5 NDCG@1 NDCG@3 NDCG@5 R@1 R@3 R@5
FLAVA Real 38.70 45.72 47.26 63.12 82.52 88.53 36.65 44.45 46.27 59.77 81.38 88.62
AI-generated 41.59 47.45 48.72 67.83 84.10 89.10 36.49 43.29 45.17 59.52 78.44 85.90
ALIGIN Real 45.43 50.62 51.65 74.10 88.38 92.44 38.13 44.71 46.33 62.20 80.52 87.00
AI-generated 43.47 49.38 50.42 70.90 87.24 91.38 36.34 43.32 45.06 59.28 78.62 85.51
BEIT-3 Real 47.45 52.15 52.87 77.40 90.32 93.20 41.24 47.16 48.63 67.25 83.72 89.52
AI-generated 45.31 50.49 51.43 73.90 88.24 91.94 38.33 44.76 46.19 62.52 80.30 85.98
We introduce three open-source widely used and powerful vision-language pre-trained models including FLAVA (Singh et al., 2022), ALIGN (Jia et al., 2021), and BEIT-3 (Wang et al., 2022) to perform retrieval on the databases that only contain real images and only contain generated images respectively. Since text-image retrieval datasets such as Flicker30k and MSCOCO are composed of real images, the training on these images will further introduce additional bias. In view of the fact that the above three models have shown excellent performance in zero-shot text-image retrieval (Wang et al., 2022), we directly use the above three models to perform retrieval under the zero-shot setting. The experimental results in Table 2 show that retrieval performance on the database containing only AI-generated images is not significantly greater than retrieval performance on the database containing only real images. It indicates that the AI-generated images in our benchmark do not introduce more visual semantics relevant to the query.
2.2.2. Human Evaluation
To further assess whether the AI-generated images in our constructed benchmark introduce additional visual semantics that are more relevant to the queries in retrieval, we conduct the human evaluation. Specifically, we invite five volunteers with master’s degrees to participate in the human evaluation. The triple consisting of the caption (i.e., the query in retrieval), real image, and its corresponding AI-generated image is the basic unit in this evaluation. We ask the five volunteers to select which image (real or AI-generated) has more relevant semantics to the caption. If the volunteers think the two images are equally relevant or irrelevant to the caption, they can select "equal". We count the proportion of selections made by volunteers on the entire benchmark and the results are shown in Table 3. Volunteers in human evaluation think that in most samples of our benchmark, real images have more or equal visual semantics relevant to the caption (i.e., the query in retrieval) than AI-generated images. This result further guarantees the fairness of our assessment of invisible relevance bias. If the AI-generated images do not have more relevant visual semantics than real images, while the text-image retrieval model still tends to rank AI-generated images higher than the real images, the invisible relevance bias does exist.
Table 3. Proportion of the selections made by volunteers in human evaluation.
Flicker30k+AIGC MSCOCO+AIGC
Which image is more relevant to the query?
Real AI Equal Real AI Equal
46.25% 13.75% 40% 45.35% 12.15% 42.5%
3. Bias Assessment
Based on the benchmark we constructed above, we can assess the impact of AI-generated images on text-image retrieval models. We not only study the impact on ranking results when AI-generated images appear in the retrieved database but also further discuss the problems caused when retrieval models are trained on the training data mixed with AI-generated images.
3.1. Text-Image Retrieval Models
From the perspective of the training data of models, our assessment includes not only the retrieval models trained from scratch on supervised text-image pairs but also vision-language models that have been pre-trained on massive image-text pairs. From the perspective of the architecture of models, our assessment includes both dual-encoder-based and fusion-based models. We do not select models that are not open source because of the relatively high implementation costs. The specific models used in the assessment include: (1) NAAF (Zhang et al., 2022) is a fusion encoder text-image retrieval model that exploits both the positive effect of matched fragments and the negative effect of mismatched fragments to jointly infer text-image similarity. (2) VSE (Chen et al., 2021) is a dual-encoder text-image retrieval model that learns to automatically adapt the best pooling strategy for visual semantic embedding. We following the papers to train NAAF and VSE on supervised text-image retrieval datasets. (3) VILT is a fusion encoder text-image matching model based on the interaction between image and text in Transformer (Vaswani et al., 2017). It is a vision-language model that has been pre-trained on massive text-image pairs. (4) FLAVA is a language and vision alignment model that learns strong representations from multimodal. It is a vision-language model pre-trained on massive text-image pairs and can be used as the dual encoder model for text-image retrieval. (5) ALIGIN is also a dual encoder vision-language that has been pre-trained on over one billion image-text pairs. (6) BEIT-3 is a general-purpose multimodal foundation model that has been pre-trained on hundreds of millions of text-image pairs and massive texts and images. It can be used as a dual-encoder text-image retrieval model.
3.2. Experimental Settings and Metrics
As the neural networks adapt not only to the features of the task signal but also to the data domain during the training process (Xu et al., 2022), our assessment is performed under the out-of-domain setting to try to mitigate potential bias introduced by the domain of the training data. Specifically, for the models that need to train from scratch on supervised text-image pairs (NAAF and VSE), we train them on dataset A and evaluate their performance on another dataset B. For the models that have been pre-trained on massive real text-image pairs and show excellent zero-shot performance in text-image retrieval, we directly use these pre-trained models to perform retrieval on the test datasets. An exception is that even though VILT has been pre-trained, it needs to be combined with a specific multi-layer perceptron to complete the text-image matching task in text-image retrieval. So we fine-tune VILT on supervised text-image retrieval datasets just like NAAF and VSE.
The main metrics in this section aim to measure the difference between the ranking of real and AI-generated images in the retrieved results. We follow (Dai et al., 2023) to measure the relative percentage difference between them as:
(2) $\rm{Relative}\triangle=\frac{2(\rm{Metric}_{real}-Metric_{AI-generated})}{Metric_{real}+Metric_{AI-generated}}\times 100\%,$
in which Metric can be the evaluation metrics for information retrieval such as NDCG@k and R@k. $\rm{Relevance\triangle}>0$ means retrieval models rank real images higher than AI-generated images, $\rm{Relevance\triangle}<0$ means retrieval models rank AI-generated images higher than real images. The absolute value of $\rm{Relevance\triangle}$ indicates the value of this bias (Dai et al., 2023).
3.3. Invisible Relevance Bias
Table 4. Performance of the retrieval models on the benchmark we constructed consisting of both real and AI-generated images. $\rm{Relevance\triangle}>0$ means retrieval models rank real images higher than AI-generated images, $\rm{Relevance\triangle}<0$ means retrieval models rank AI-generated images higher than real images. The absolute value of $\rm{Relevance\triangle}$ indicates the degree of the bias.
Flicker30k+AIGC MSCOCO+AIGC
NDCG@1 NDCG@3 NDCG@5 R@1 R@3 R@5 NDCG@1 NDCG@3 NDCG@5 R@1 R@3 R@5
Models trained from scratch
[…]
AI-generated 19.59 29.68 31.86 31.96 59.78 68.34 13.56 20.93 23.37 22.12 43.21 53.90
$\rm{Relative}\triangle$17.81 -9.00 -8.05 -17.81 -5.8 -4.36 -13.53 -3.64 -2.22 -13.53 -1.29 -1.24
[…]
AI-generated 17.04 26.04 28.31 27.79 52.70 61.70 10.75 17.87 20.33 17.54 37.50 47.24
$\rm{Relative}\triangle$23.57 -10.63 -7.86 -23.57 -6.45 -2.31 -1.13 -0.73 0.62 -1.13 -0.66 1.63
Pre-trained Vision-Language Models
[…]
AI-generated 37.61 44.86 46.36 61.33 81.34 87.26 27.01 36.81 38.87 44.06 70.99 79.12
$\rm{Relative}\triangle$148.85 -83.78 -72.44 -148.85 -58.32 -40.69 -72.81 -34.49 -29.00 -72.81 -21.36 -13.21
[…]
AI-generated 25.48 39.10 40.91 41.56 78.38 85.44 21.31 33.23 35.49 34.76 67.24 76.16
$\rm{Relative}\triangle$14.6 -4.95 -4.59 -14.6 -1.93 -1.49 -12.41 -5.65 -4.63 -12.41 -3.48 -1.88
[…]
AI-generated 24.40 39.54 41.12 39.80 80.50 86.68 21.24 34.55 36.63 34.64 70.86 79.08
$\rm{Relative}\triangle$0.72 -2.17 -1.41 -0.72 -2.97 -1.44 0.62 -3.90 -3.01 0.62 -5.50 -3.72
[…]
AI-generated 20.04 30.43 32.71 32.70 61.30 70.30 18.29 31.21 33.50 29.85 63.30 72.30
$\rm{Relative}\triangle$13.38 -2.69 -1.69 -13.38 0.97 2.25 -11.51 -4.90 -4.32 -11.51 -0.32 0.28
The performance of text-image retrieval models on the benchmark we constructed consisting of both real and AI-generated images are shown in Table 4. Overall, invisible relevance bias widely exists in text-image retrieval models, that is, text-image retrieval models tend to rank AI-generated images higher than real images even though they have very similar semantics. Specifically, the following conclusions can be made: (1) The invisible relevance bias exists in both the models trained from scratch and the vision-language models that have been pre-trained on massive supervised text-image pairs. (2) The invisible relevance bias exists in both dual-encoder-based and fusion-encoder-based retrieval models. (3) The invisible relevance bias has a relatively greater impact on the Top-1 retrieved image. In the retrieved list, the Top-1 item is most likely to be clicked by users, which means that invisible relevance bias introduced by AI-generated images will have a huge impact on users’ actual search and click results.
3.4. More Serious Bias Caused by Training
(a) $\rm{Relative}\triangle$ on R@k of Flicker30k
(b) $\rm{Relative}\triangle$ on NDCG@k of Flicker30k
(c) NDCG@k on only real or AI-generated images of Flicker30k
(d) $\rm{Relative}\triangle$ on R@k of MSCOCO
(e) $\rm{Relative}\triangle$ on NDCG@k of MSCOCO
(f) NDCG@k on only real or AI-generated images of MSCOCO
Figure 3. In-domain and out-of-domain assessment results of VSE trained on the datasets that reconstructed based on the training set of Flicker30k. We change the ratio of AI-generated images in the datasets (X-axis) while keeping the total number of samples in the training set unchanged. The model is tested on the test set of Flicker30k+AIGC (in-domain) and MSCOCO+AIGC (out-of-domain) respectively that we constructed in Section 2.
(a) $0\%$ on Flicker30k
(b) $20\%$ on Flicker30k
(c) $40\%$ on Flicker30k
(d) $60\%$ on Flicker30k
(e) $80\%$ on Flicker30k
(f) $100\%$ on Flicker30k
(g) $0\%$ on MSCOCO
(h) $20\%$ on MSCOCO
(i) $40\%$ on MSCOCO
(j) $60\%$ on MSCOCO
(k) $80\%$ on MSCOCO
(l) $100\%$ on MSCOCO
Figure 4. Distribution of the relevance scores between captions and images estimated by retrieval models trained on the datasets mixed with different ratios of AI-generated images. Flicker30k is an in-domain setting and MSCOCO is an out-of-domain setting.
An essential point that needs to be noticed is that AI-generated images do not only have an impact on the inference stage of the retrieval models. Due to their large-scale distribution on the Internet and the bias that has been found in Section 3.3, AI-generated images are very likely to be mixed into the training data of retrieval models in the future. In this section, we delve deeper into the repercussions on retrieval performance and invisible relevance bias when AI-generated images are mixed into the training of retrieval models. Specifically, we explore the impact on the retrieval models when the training set is mixed with generated images by changing the ratio between real and generated images in the supervised text-image training set. To ensure a comprehensive and accurate assessment, our experiments focus on the model trained from scratch (VSE). This is because pre-trained vision-language models have been trained on a large number of real text-image pairs, which may introduce biases in evaluating the impact of different ratios of generated images in the training data. In this experiment, we reconstruct the training set of Flicker30k by replacing a certain proportion of real images with AI-generated images. Specifically, for each real image ($I_{r}$) in the training set of Flicker30k, we use the method in Section 2 to generate its corresponding AI-generated image ($I_{g}$). Then, we introduce the ratio $\alpha$, which means that in our reconstructed training data, the paired images for $\alpha$ percentage of captions are AI-generated images, and for $(100-\alpha)$ percentage of captions are real images. We change the ratio while keeping the total number of samples in the training set unchanged. We evaluate the performance of the trained model in the test set of Flicker30k+AIGC (in-domain setting) and MSCOCO+AIGC (out-of-domain setting).
The experimental results are shown in Figure 3 and Figure 4. Figure 3 illustrates that as the ratio increases, the ranking disparity between generated images and real images widens, with the retrieval model exhibiting a greater inclination to rank generated images higher. Meanwhile, Figure 4 demonstrates that with an increasing ratio, the mean of the score distribution for generated images rises. Simultaneously, the difference in score distribution between generated images and real images grows, and the scores of AI-generated images gradually become greater and greater than those of real images. The following conclusions can be drawn from these results: (1) Retrieval models trained on the datasets mixed with AI-generated images exhibit more serious invisible relevance bias in both in-domain and out-of-domain settings. (2) The invisible relevance bias tends to become more obvious as the ratio of AI-generated images in the training datasets increases in both in-domain and out-of-domain settings. (3) As the ratio of AI-generated images in the training datasets increases, the retrieval performance on real images is gradually damaged. This phenomenon is more obvious in the in-domain setting.
The experimental results in this section also reveal a vicious cycle of falling into invisible relevance bias. Specifically, the invisible relevance bias of text-image retrieval models causes the AI-generated images to have a higher chance of being obtained from massive data, which makes them more likely to be mixed into the training of retrieval models, leading to more serious bias and forming a vicious cycle. This ultimately results in users’ search results being surrounded by AI-generated images. This indicates that attention should be paid to alleviating the ranking bias caused by AI-generated images.
4. Our Method in debiasing
Section 3 reveals that invisible relevance bias widely exists in different text-image retrieval models and the training on data mixed with AI-generated images makes this bias more serious, which forms a vicious cycle and results in users’ search results being surrounded by AI-generated images. Therefore, it is very important to propose a method to mitigate this bias. In this section, we propose an effective method to achieve it by measuring and optimizing the invisible relevance bias in training. The specific details of our method will be introduced in the following contents.
4.1. Design of Debiasing Method
Our method alleviates the retrieval model’s preference bias for generated images by measuring and optimizing the additionally introduced invisible relevance in the training. For a caption-image pair $(C_{i},I^{r}_{i})$ in the training set, whether the dual-encoder or fusion-encoder architecture text-image retrieval models, the training process can be summarized as estimating the relevance score between $C_{i}$ and $I^{r}_{i}$, and using contrastive loss or regression loss as the optimization objective, which can be described as:
$s$ $=R(C_{i},I^{r}_{i};\theta),$
$\theta_{\text{optimal}}$ $=\arg\min_{\theta}\mathcal{L}(s,y;\theta),$
in which $R(\cdot;\theta)$ is the relevance scoring function of the retrieval model such as cosine similarity between embeddings in the dual-encoder model and neural networks in the fusion-encoder model, $\theta$ is the set of parameters of the model, $\mathcal{L}$ is the loss function such as contrastive loss or regression loss, $y$ is the label. In our method, for each real image $I^{r}_{i}$ in the training data, we use the method in Section 2 to generate its corresponding AI-generated image $I^{g}_{i}$. Then we can get the caption-real-AI triple as $(C_{i},I^{r}_{i},I^{g}_{i})$, in which $C_{i}$ is the paired caption for the real image $I^{r}_{i}$. We introduce the contrastive loss to get the difference in relevance scores between $I^{r}_{i}$ and $I^{g}_{i}$ with respect to the caption $C_{i}$ as:
$\triangle s(I^{g}_{i},I^{r}_{i})$ $=R(C_{i},I_{i}^{g};\theta)-R(C_{i},I^{r}_{i};\theta),$
which can measure the additional invisible relevance introduced by the AI-generated image $I^{g}_{i}$ for $C_{i}$ compared with the real image $I^{r}_{i}$. Therefore, this can be used as a part of the optimization objective to mitigate the invisible relevance bias. It is because minimizing the difference between $R(C_{i},I_{i}^{g};\theta)$ and $R(C_{i},I^{r}_{i};\theta)$ in training can make the retrieval model eliminate the additional score estimated for $I^{g}_{i}$. In all triples in the training data, we consider the caption-real-AI triples $(C_{i},I^{r}_{i},I^{g}_{i})$ whose $\triangle s(I^{r}_{i},I^{g}_{i})$ is greater than 00 and perform Bernoulli sampling from these triples with probability $\beta$ to get the target triple set $\mathbb{B}$ for debiasing optimization as:
$\mathbb{B}=\{(C_{i},I^{r}_{i},I^{g}_{i})\mid(C_{i},I^{r}_{i},I^{g}_{i})\sim\text{Bernoulli}(\beta)\cdot\mathbb{I}(\triangle s(I^{g}_{i},I^{r}_{i})>0)\}.$
The reason why we only sample triples with probability $\beta$ is to adjust the tolerance of the retrieval models to the AI-generated images. The higher the probability $\beta$, the more likely the retrieval models are to rank the AI-generated images to a lower position. The total optimization objective in training is:
$\theta_{\text{optimal}}=\arg\min_{\theta}(\sum_{s_{i},y_{i}\in\mathbb{A}}\mathcal{L}(s_{i},y_{i};\theta)+\sum_{I^{r}_{i},I^{g}_{i}\in\mathbb{B}}\triangle s(I^{g}_{i},I^{r}_{i})),$
in which $\mathbb{A}$ is the set of all samples in the training data.
4.2. Evaluation of Debiasing
Table 5. Performance of the retrieval models on the benchmark we constructed consisting of both real and AI-generated images with different sampling probability $\beta$ in our debiasing method. $\rm{Relevance\triangle}>0$ means retrieval models rank real images higher than AI-generated images, $\rm{Relevance\triangle}<0$ means retrieval models rank AI-generated images higher than real images. When $\rm{Relevance\triangle}<0$, the absolute value of $\rm{Relevance\triangle}$ indicates the value of this bias.
Flicker30k+AIGC (In-domain) MSCOCO+AIGC (Out-of-domain)
$\beta$ $\beta$
w/o debias 50% 60% 70% 80% 90% 100% w/o debias 50% 60% 70% 80% 90% 100%
$\rm{Relative}\triangle$ on NDCG@1 -10.35 -1.406 31.42 62.77 91.71 112.06 129.20 -13.53 -1.384 45.35 80.67 114.63 140.23 154.43
$\rm{Relative}\triangle$ on NDCG@3 -4.31 -0.656 15.08 32.85 50.08 65.44 77.31 -3.64 -0.354 27.61 53.11 78.92 101.67 114.31
$\rm{Relative}\triangle$ on NDCG@5 -4.37 -0.876 13.13 27.84 42.28 55.68 65.31 -2.22 -0.214 23.47 46.42 69.62 90.28 102.89
NDCG@1 on only real images 30.57 33.44 33.15 33.26 33.12 33.09 33.20 18.50 21.09 21.48 21.32 20.52 20.43 20.01
NDCG@1 on only real images 37.95 40.44 40.32 40.53 40.38 40.13 40.31 25.66 28.92 29.05 28.78 28.32 28.05 27.65
NDCG@5 on only real images 39.78 42.29 41.98 42.18 42.05 41.93 42.10 28.28 31.55 31.69 31.43 30.97 30.56 30.02
(a) $\beta=50\%$ on Flicker30k
(b) $\beta=60\%$ on Flicker30k
(c) $\beta=70\%$ on Flicker30k
(d) $\beta=80\%$ on Flicker30k
(e) $\beta=90\%$ on Flicker30k
(f) $\beta=100\%$ on Flicker30k
(g) $\beta=50\%$ on MSCOCO
(h) $\beta=60\%$ on MSCOCO
(i) $\beta=70\%$ on MSCOCO
(j) $\beta=80\%$ on MSCOCO
(k) $\beta=90\%$ on MSCOCO
(l) $\beta=100\%$ on MSCOCO
Figure 5. Distribution of the relevance scores between captions and images estimated by retrieval models with different sampling probability $\beta$ in our debiasing method. Flicker30k is an in-domain setting and MSCOCO is an out-of-domain setting.
(a) $\beta=50\%$ on Flicker30k
(b) $\beta=60\%$ on Flicker30k
(c) $\beta=70\%$ on Flicker30k
(d) $\beta=80\%$ on Flicker30k
(e) $\beta=90\%$ on Flicker30k
(f) $\beta=100\%$ on Flicker30k
(g) $\beta=50\%$ on MSCOCO
(h) $\beta=60\%$ on MSCOCO
(i) $\beta=70\%$ on MSCOCO
(j) $\beta=80\%$ on MSCOCO
(k) $\beta=90\%$ on MSCOCO
(l) $\beta=100\%$ on MSCOCO
Figure 6. T-SNE visualization of the embeddings of real and AI-generated images obtained from the image encoder in VSE.
We evaluate the effect of our debiasing method from three perspectives: (1) How our method affects the ranking difference between real images and AI-generated images. (2) How our method affects the distribution of relevance scores between captions and images. (3) How our method affects the representations of real and AI-generated images encoded by the image encoder of the retrieval model. All of these three evaluations are performed in both in-domain and out-of-domain settings, which can provide a comprehensive evaluation of our proposed method.
Table 5 shows how the $\rm{Relative\triangle}$ and retrieval performance changes with the sampling probability $\beta$. The results indicate that: (1) Our method not only effectively alleviates the retrieval model’s preference for AI-generated images, but also makes real images ranked significantly higher than AI-generated images in the final ranked list. (2) As the sampling probability $\beta$ increases, the retrieval model can better avoid the interference of AI-generated images and better expose real images to users. (3) When $\beta$ is $0.5$, retrieval models can achieve a fair ranking between real and generated images with little bias. (4) Our method not only does not impair the retrieval performance but also enhances it. It is mainly because, under the constraints of our method, AI-generated images can be seen as the hard negatives in training, which enhances the ability of the retrieval model to estimate the relevance, especially for distinguishing the images with very similar semantics.
Figure 5 shows the distribution of the relevance scores between captions and images estimated by retrieval models with different sampling probability $\beta$. The results indicate that our method effectively reduces the relevance scores to AI-generated images estimated by the retrieval model. With an increase in the sampling probability ($\beta$), the disparity in score distribution between the generated images and the real images expands, and the scores of real images gradually become greater and greater than those of AI-generated images.
Figure 6 shows the T-SNE visualization of the embeddings of real and AI-generated images acquired from the image encoder in VSE. The findings reveal that, with an increase in the sampling probability ($\beta$), the representations of generated and real images become more distinguishable. This observation indicates that our method empowers the retrieval model to capture the feature differences between generated and real images, reflecting these distinctions in the representations.
5. Causes of Invisible Relevance Bias
In this section, we use the debiased model proposed in Section 4 to reversely analyze the causes of the invisible relevance bias. Specifically, we obtain the transformations made by the debiased retriever on AI-generated images by comparing the difference in image representations between the original retriever and the debiased retriever. The reverse process of these transformations can be seen as the cause of invisible relevance bias. We also design the relevant experiments to support this point of view. Details will be introduced in the following contents.
5.1. Transformations in Debiasing
To simplify the analysis process, we focus on the dual-encoder retrieval model that estimates the relevance score by dot product or cosine similarity between text and image representations. For the same text, image representation can directly affect the estimation of the relevance score, which can help us to find the most direct causes of invisible relevance bias, therefore, image representation is the main object in our analysis. Specifically, we aim to analyze the difference between image representations encoded by the original retriever and debiased retriever to get the transformations of debiasing from this difference and reverse these transformations to explore the causes of the invisible relevance bias.
Given the image encoder $v(\cdot;\theta)$ of original text-image retrieval model and image encoder $v^{d}(\cdot;\theta^{\prime})$ of our proposed debiased retrieval model, a AI-generated image $I^{g}$, we can get the representations of $I^{g}$ encoded by $v(\cdot;\theta)$ and $v^{d}(\cdot;\theta^{\prime})$ respectively as:
$\bm{r}$ $=v(I^{g};\theta),$
$\bm{r^{d}}$ $=v^{d}(I^{g};\theta^{\prime}).$
For the normalized text representation $\bm{t}=[t_{1},t_{2},...,t_{n}]$, image representation $\bm{r}=[r_{1},r_{2},...,r_{n}]$ from the original retriever and image representation $\bm{r^{d}}=[r^{d}_{1},r^{d}_{2},...,r^{d}_{n}]$ from the debiased retriever, the relevance $score(\bm{t},\bm{r})$ and $score(\bm{t},\bm{r^{d}})$ can be calculated as:
$score(\bm{t},\bm{r})$ $={t_{1}}{r_{1}}+{t_{2}}{r_{2}}+{t_{3}}{r_{3}}+...+{t_{n}}{r_{n}},$
$score(\bm{t},\bm{r^{d}})$ $={t_{1}}{r^{d}_{1}}+{t_{2}}{r^{d}_{2}}+{t_{3}}{r^{d}_{3}}+...+{t_{n}}{r^{d}_{n}}.$
Therefore, the adjustment of relevance score between text $t$ and AI-generated image $I^{g}$ in the debiased model is essentially changing the values of each element ($r_{i}$) in $\bm{r}$, which can be described as:
$score(\bm{t},\bm{r^{d}})$ $={t_{1}}{r^{d}_{1}}+{t_{2}}{r^{d}_{2}}+...+{t_{n}}{r^{d}_{n}}$
$={t_{1}}{(r_{1}+\triangle r_{1})}+{t_{2}}{(r_{2}+\triangle r_{2})}+...+{t_{n}}{(r_{n}+\triangle r_{n})}.$
The transformations $\bm{p}$ in the debiased model can be represented by a vector with the same dimensions as $r$ and $r^{d}$:
$\bm{p}$ $=[\triangle r_{1},\triangle r_{2},...,\triangle r_{n}]$
$=[p_{1},p_{2},...,p_{n}]$
$=r\ominus r^{d},$
in which $\ominus$ means bitwise subtraction between $r$ and $r^{d}$. Then, we perform two-dimensional visualization of the $\bm{r}$, $\bm{r^{d}}$ and $\bm{p}$ of all images in datasets to try to find the patterns from them. The T-SNE visualization is shown in Figure 7. This result shows that compared with the scattered image representations, the transformations vector $\bm{p}$ shows an obvious aggregation phenomenon. This phenomenon shows that there is consistency in the transformations performed by the debiased retriever on AI-generated images with very different semantics. The role of the debiased retriever is stripping the transformations vector $\bm{p}$ out of the representations of AI-generated images.
Figure 7. T-SNE visualization of image representations and transformations in datasets.
5.2. Reversing the Transformations
Debiased retriever modifies each element ($r_{i}$) of the representation $\bm{r}$ from the original retriever according to the value of the corresponding element $p_{i}$ in transformations vector $\bm{p}$ and gets the debiased representation $\bm{r^{d}}$, which can be described as:
$r^{d}_{i}=r_{i}+p_{i},r_{i}\in\bm{r},p_{i}\in\bm{p},r^{d}_{i}\in\bm{r^{d}}.$
We can reverse this process to get the causes of the invisible relevance bias. This is because the transformation made by a debiased retriever to make a biased AI-generated image become unbiased is exactly the difference between the biased AI-generated images and the real images. That is, the reason why an unbiased representation $\bm{r^{d}}$ becomes the representation $\bm{r}$ with bias is that the reverse transformation ($-p_{i}$) is done on $r^{d}_{i}$, which can be described as:
$r_{i}=r^{d}_{i}-p_{i}.$
Therefore, we conclude that an unbiased representaion $\bm{r^{d}}$ becomes the representaion $\bm{r}$ with bias because $\bm{r^{d}}$ is added bitwisely by a vector $\bm{-p}=[-p_{1},-p_{2},-p_{3},...,-p_{n}]$. Combining the phenomenon shown in Figure 7 that there is consistency in the transformations vector $\bm{p}$ on AI-generated images with very different semantics, we can get the causes of invisible relevance bias in AI-generated images: AI-generated images cause the image encoder in the retriever to embed additional information to their representations. This information has the following three characteristics:
(1) When this information is embedded into the representation outputted from the visual encoder, it can amplify the relevant information to get the higher relevance score.
(2) This information cannot be reflected in visual semantics in a visible way, but can only be extracted by neural network-based visual models.
(3) This information has a certain consistency in AI-generated images with different semantics. This information may be like the watermark that is universal information for the image generation model and can be expressed by different neural network-based visual models such as image encoder in retriever.
For the above three points, we propose experiments to support them, which will be introduced in the following contents.
5.3. Supporting Experiments
This section aims to use experimental results to support three characteristics in the causes of invisible relevance bias mentioned above.
For the first point, we design a direct experiment to support it. We apply the bitwise addition of the reverse transformations vector $\bm{-p}$ to the representations of real images encoded by the original, non-debiased retriever and detect whether the bias can be eliminated. The intention for this is that if this additional information (transformations vector $\bm{-p}$ ) is indeed the cause of the AI-generated images being ranked higher, then when we add these transformations to the representation of real images, the real images will also be ranked higher, thereby eliminating the ranking difference between the real and generated images. The experimental results are shown in Table 6. It is very surprising that the ranking advantage of generated images over real images caused by invisible relevance bias is not only eliminated but reversed by simply bitwisely adding the transformations vector $\bm{-p}$ to the representation of the real images without any debiasing training. This proves that the reverse transformations vector $\bm{-p}$ we found is an important cause of the invisible relevance bias and it is implicit in the AI-generated images and can be embedded into the representations by the image encoder.
For the second point, human evaluation in Section 2.2.2 and retrieval performance in Section 2.2.1 have shown that AI-generated images do not introduce additional visual semantics compared to their real images, indicating that the additional information is invisible. Besides, the ranking bias detected in Section 3.3 and Section 3.4 shows that this additional information can be captured by the image encoder and reflected in the relevance score between the queries.
For the third point, T-SNE visualization of image representations and transformations vector $\bm{p}$ in Figure 7 has shown that compared with the scattered image representations, the transformations vector $\bm{p}$ show an obvious aggregation phenomenon. This proves that for AI-generated images with different semantics, the debiased model only needs roughly consistent transformations on representations to remove the bias, which means that there is a certain consistency in the additional information for AI-generated images encoded by the image encoder.
Table 6. Effect of reverse transformations vector $\bm{-p}$ on invisible relevance bias. The retrieval model is VSE trained on Flicker30k without any debiasing training.
Flicker30k+AIGC
Relative $\triangle$ on
NDCG@1 NDCG@3 NDCG@5 R@1 R@3 R@5
Original -10.35 -4.31 -4.37 -10.35 -4.72 -4.06
Add $\bm{-p}$ to Real 17.85 4.54 2.99 17.85 -0.28 -1.17
6. Related Work
6.1. Artificial Intelligence Generated Content
With the development of generation models, AI-generated content (AIGC) becomes more realistic and less discernible (Agnese et al., 2020; Cao et al., 2023; Wu et al., 2023). The use of AI to generate content has greatly changed the way of content generation. It reduced the cost of content generation and improved the efficiency than traditional compared to traditional content generation pipeline biased on humans (Hanley and Durumeric, 2023; Spitale et al., 2023). For example, Stable Diffusion (Yang et al., 2023) and DALL-E-3 (Betker et al., 2023) can generate high-quality images by simply following a short text description of uses’ input. However, as AIGC is gradually applied to various content production pipelines in society and the Internet, some potential risk concerns are also worthy of consideration. Some studies have found that AI-generated content has risks in many aspects, such as discrimination, privacy leakage, ethics, and safety (Deshpande et al., 2023; Zhuo et al., 2023; Chen and Shu, 2023; Jiang et al., 2023; Su et al., 2023). A recent study suggests that texts generated by AI will introduce bias in text modal information retrieval, causing the retriever to rank them higher (Dai et al., 2023). Therefore, in the context of AIGC being widely used, the potential risks it may bring should be carefully studied.
6.2. Bias and Fairness in Information Retrieval
Due to the influence of training data, indexes, models, etc., the ranking results of information retrieval will produce certain deviations (Gao and Shah, 2021). These biases are specifically reflected in various aspects such as gender (Azzopardi, 2021), cognition (Bigdeli et al., 2021), understandability (Zuccon, 2016), and retrievability (Wilkie and Azzopardi, 2014). Kulshrestha et al. (Kulshrestha et al., 2017) formulated a framework for quantifying bias, addressing its origins in both the data source and the ranking system. Beutel et al. (Beutel et al., 2019), Kuhlman et al. (Kuhlman et al., 2019), and Yao and Huang (Yao and Huang, 2017) introduced pairwise comparisons focusing on utility and prediction errors. Geyik et al. (Geyik et al., 2019) and Yang along with Stoyanovich (Yang and Stoyanovich, 2017) suggested fairness measures based on distances for ranked outputs, grounded in statistical parity. Gao (Gao, 2021) advocated for a unified evaluation metric tailored for fairness-aware ranking algorithms. Diaz et al. (Diaz et al., 2020) proposed metrics specifically designed for assessing fairness in stochastic rankings. Numerous studies have conducted comparisons among existing fairness metrics (Geyik et al., 2019; Diaz et al., 2020; Raj and Ekstrand, 2020; Sapiezynski et al., 2019). In this paper, we focus on the ranking bias caused by AI-generated Images in text-image retrieval.
7. Conclusion
This paper delves into the exploration of source bias in text-image retrieval, focusing particularly on the phenomenon of invisible relevance bias introduced by AI-generated images. Through the construction of a comprehensive benchmark with visually similar semantics between real and AI-generated images, we unveil the prevalence of this bias in text-image retrieval models. Our findings underscore that AI-generated images tend to be ranked higher by retrieval models, despite lacking more visually relevant features to the query than real images. Regardless of the model’s training data and architecture, ranging from models trained from scratch to those pre-trained on extensive image-text pairs, and encompassing both dual-encoder and fusion-encoder models, the invisible relevance bias persists. Moreover, mixing AI-generated images into the training data exacerbates the invisible relevance bias, creating a self-perpetuating vicious cycle where these AI-generated images gain more exposure in massive datasets, increasing their likelihood of being integrated into retrieval model training, thus intensifying the bias further. In response to these challenges, we propose a novel debiasing training method to mitigate the invisible relevance bias effectively. This method can dynamically adjust the preference to AI-generated images of text-image retrieval models by adjusting the sampling probability. Then, we use our proposed debiasing method to reversely determine that the causes of invisible relevance is that the AI-generated images cause the image encoder to introduce additional information into their representation. This information has a certain consistency in different generated images and can make the retriever estimate a higher relevance score. Our proposed method and view, supported by rigorous experiments, serves as a crucial step toward understanding and rectifying the invisible relevance bias cased by AI-generated images in text-image retrieval models.